OpenStudy (anonymous):

what is the perimeter of a right triangle in which th longer leg is 4 units longer than the hypotenuse ?

OpenStudy (anonymous):

the thing has to add up to one 80 and the hypotneuse has to be bigger than both sides do have multiple choice so i can explain better?

OpenStudy (anonymous):

u do need help still right @veunta&NuNu

OpenStudy (anonymous):

a:) 12units b:)16units C:)20 units D:)48 units

OpenStudy (anonymous):

do you know which are the another side besides 4 is? and i meant 180 in my first reply

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do u know the formular \[c ^{2}=a ^{2}+b ^{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok c^2=4+b^2, it has to eaqul 360 from being squared and then add regular a+b+c to get the perimiter understand?

OpenStudy (anonymous):

yeah i kinda understand ,,,is it anyway other way you can helpp me to find the answer

OpenStudy (anonymous):

text the theory \[c ^{2}=4^{2}+b ^{2}\]

OpenStudy (anonymous):

4*4=16 so c^2=16+b^2

OpenStudy (anonymous):

that is the star of your perimitor but lets show you how to get 360 first

OpenStudy (anonymous):

4+_=360

OpenStudy (anonymous):

that would be what?

OpenStudy (anonymous):

56

OpenStudy (anonymous):

now split 56

OpenStudy (anonymous):

28

OpenStudy (anonymous):

28 and what?

OpenStudy (anonymous):

then what ?

OpenStudy (anonymous):

ok so 4^2+28^2=28^2 is what

OpenStudy (anonymous):

here u want me to show u a diff way?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Let the longer leg be X ok? Then the snorter leg must be X-4 Now the Pythagorean Theorem tells us that in right-angled triangles, (short leg)^2 + (longer leg)^2 =hypotenuse^2 (X-4)^2 +( X )^2 = 6^2 X^2 -8X + 16 + X^2 =36 2X^2-8X-20 =0 X^2 -4X -10=0 understand so far?

OpenStudy (anonymous):

should i keep going?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

At this point I can't factor by eye-balling it, so I'll have to use the quadratic formula That formula says X= {-b+ or - sq rt(b^2-4ac)}/2a' where a =X^2 coefficient, in this case +1 b=X coefficient, in this case -4 c == constant term, in this case -10 X= {-(-4)+ or - sqrt(-4^2 -4(1)(-10)}/2 X ={4 + or - sqrt(16+40)}/2 X ={4 + or - sqrt(56)}/2 X={4 + or -sqrt4(sqrt14)}/2 X={4 + or - 2rt14}/2 X=2{2+ or -rt14}/2 X= 2+ rt14 or 2 -rt14 rt14= 3.74 So, X = 5.74 or -1.74 Clearly -1.74 is out. Therefore the side X is 5.74and since this is the longer leg, the shorter leg must be 5.74-4=1.74 Let's check our work. (5.74)^2 + (1.74)^2 =36 32.95 +3.03=36 35.98 = 36

OpenStudy (anonymous):

i rounded that to 48 understand?

OpenStudy (anonymous):

yes i understand can you hel[p me with the next qestion

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

if the length of the hypotenuse ofan isosceles right triangle is 12cm,what is the length of each of the other two sides ?

OpenStudy (anonymous):

Use Pythagoras' Theorem a² + b² = 12² The triangle is isosceles so a=b 2a² = 12² a² = 72 a = 6√2 ≈ 8.49cm The lengths of each of the other two sides are about 8.49cm You could also use the fact that the angles of the triangle would be 90°, 45° and 45° and use the sin and cos values of 45° (sin45° = cos45° = 1/√2) to arrive at the same result. understand???

OpenStudy (anonymous):

answer would be ?

OpenStudy (anonymous):

about 8.49

OpenStudy (anonymous):

answer choices are 5^2 , 6 cm ,6^2 cm ,or 7 cm

OpenStudy (anonymous):

then 7, rounding

OpenStudy (anonymous):

if the length of the hypotenuse of an isosceles right triangle is 14cm,what is the lenght of the other two sides

OpenStudy (anonymous):

come on bud its the same thing lol you try it and i'll help, show me what we do first